OpenStudy (anonymous):

Simplify the expression. 3^7/3^5 A. 3^35 B. 3^12 C. 1/3^9 D. 9

OpenStudy (lgbasallote):

\[\large \frac{3^7}{3^5} = 3^{7-5}\]

OpenStudy (anonymous):

D ?

sam (.sam.):

\[\large \frac{3^7}{3^5} = 3^{7-5} \] \[\large 3^{7-5} = 3^{2} \] \[\large 3^{2} =?\]

OpenStudy (anonymous):

\[3^{7} \div 3^{5} =\] dividing with same base, just subtract the exponents \[7 - 5\] whih is equal to 2 the answer would be \[3^{2}\] or 9